### v1

# Condensed Sample-Guided Model Inversion for Knowledge Distillation

Abstract
Knowledge distillation (KD) is a key element in neural network compression that allows knowledge transfer from a pre-trained teacher model to a more compact student model. KD relies on access to the training dataset, which may not always be fully available due to privacy concerns or logistical issues related to the size of the data. To address this, ”data-free” KD methods use synthetic data, generated through model inversion, to mimic the target data distribution. However, conventional model inversion methods are not designed to utilize supplementary information from the target dataset, and thus, cannot leverage it to improve performance, even when it is available. In this paper, we consider condensed samples, as a form of supplementary information, and introduce a method for using them to better approximate the target data distribution, thereby enhancing the KD performance. Our approach is versatile, evidenced by improvements of up to $11.4\%$ in KD accuracy across various datasets and model inversion-based methods. Importantly, it remains effective even when using as few as one condensed sample per class, and can also enhance performance in few-shot scenarios where only limited real data samples are available.
Introduction
Knowledge distillation (KD) (Hinton, Vinyals, and Dean 2015) is an effective technique for reducing the resource demand of large pre-trained models. It functions by transferring the task knowledge of a large pre-trained model called the “teacher” to a more compact “student” architecture. One restrictive aspect of KD is its dependence on the training dataset of the target task. Such data might not be available in full scale, making distillation ineffective or even infeasible. For example, the training data could be privacy sensitive or large in memory size, aggravating its relocation. Addressing this issue, several works examine methods to perform knowledge distillation even when no sample from the training set is available at the compression stage. These are called data-free distillation techniques and are mainly based on generating synthetic samples to substitute the real training set (Lopes, Fenu, and Starner 2017; Yu et al. 2023; Liu et al. 2024). To generate informative samples that are aligned with the real data distribution, the information encapsulated in the teacher model is leveraged, which is called model inversion.
Figure 1: Illustration of our motivation for using condensed samples as prototypes for synthetic data.
In certain situations, some information from the target dataset can be communicated. However, as model inversion technique is primarily designed to facilitate knowledge distillation when no training data is available, even when such information is supplied, existing methods cannot leverage it for enhancing KD performance. In this work, we explore the potential of condensed samples (Zhao, Mopuri, and Bilen 2021) from target data as a form in which supplementary information can be available, and introduce a method to refine the synthetic data generation process in model inversion whenever they are available. Our method uses condensed sampled as prototypes to propagate new samples from the training distribution as illustrated in Figure 1. Condensed samples are synthetic samples that summarize large batches from target dataset. They provide certain privacy benefits, as studied in (Dong, Zhao, and Lyu 2022) and can be produced with modest memory and time resources through recent methods (Zhou, Nezhadarya, and Ba 2022; Zhao and Bilen 2023; Feng, Vedantam, and Kempe 2023). These qualities render condensed samples suitable for scenarios in which the privacy considerations prohibit exposure of individual training samples.
A trivial approach is simply combining the condensed samples with those obtained by model inversion to distill the student. However, preliminary experiments suggest that this does not improve accuracy, which is likely caused by the domain gap between synthetic and condensed samples. With this observation, we propose using them to refine the quality of synthetic samples generated by model inversion instead. Our approach mainly targets improving the assumptions made by model inversion methods about the target data distribution. One such assumption is that real samples will be classified by the pre-trained teacher with high confidence, receiving a one-hot-like prediction vector. However, in many cases, models can give less confident predictions to real samples, yet still classify them correctly. Thus, guiding the generation process based on such assumptions on the confidence of teacher’s predictions can produce synthetic samples that do not follow the real training distribution.
Our method utilizes a small set of condensed samples to query the teacher model and extract more realistic impressions related to the target dataset. Specifically, these samples are fed to the teacher model for estimating the per-class feature distributions of the target data. Then, the model inversion process is conditioned to produce synthetic samples following a similar feature distribution as the condensed ones. This is achieved by configuring a feature discriminator that competes against the sample generator to distinguish condensed samples from the ones generated by model inversion. By doing so, the generator is forced to produce synthetic samples with semantic similarity to the condensed ones, thus reducing the risk of a domain gap and improving accuracy. One of the major advantages of our approach is its versatility, as it can be applied on top of any model inversion method to improve the accuracy of the student. This is evident from the experimental evaluation, where we record accuracy improvements of up to $11.44\%$ compared to different state-of-the-art KD baselines across multiple model pairs and datasets. The effectiveness of our method is more pronounced for teacher-student pairs with little structural similarity. In addition, remarkably, even using as few as one condensed sample per class results in a noticeable accuracy improvement. Moreover, our method is also applicable to scenarios where a limited amount of real samples from the training set are accessible, such as in few-shot learning scenarios. Experimental results exhibit the advantage of using synthetic data samples generated with guidance from the real samples against pure few-shot KD methods.
In other words, if condensed samples are available, our method can use them to enhance the performance of model inversion-based KD (data-free KD). Alternatively, it can also enhance KD performance if only a few samples from the real dataset are available. Otherwise, it simply reduces to any data-free KD method that it is coupled with. Lasty, our method does not introduce a significant time overhead, which is limited to the training cost of our lightweight discriminator. Our contributions in this work can be summarized as follows.
  • A novel approach to take advantage of the supplementary information that can be available from target data, which can be plugged into any model inversion method.
  • Investigating the applicability of our approach where either condensed samples or a few real samples from target data can be provided.
Related Work
Knowledge Distillation (KD)
KD (Hinton, Vinyals, and Dean 2015) trains a compact “student” neural network model to approximate the decision space of a more complex one called the “teacher”. The inclusion of the soft guidance supplied by the teacher enriches the limited information the student receives from the one-hot encoded class labels. As a result, the student can achieve better performance than supervised training alone can provide. While most commonly, the logit scores or the softmax probabilities of the teacher are considered for regularization (Romero et al. 2014), activation maps or attention scores can also be used (Zagoruyko and Komodakis 2016a).
Few-Shot KD
In few-shot KD, only a small subset of real training samples are accessible. To avoid over-fitting, available methods typically reduce the number of parameters required to train the student network. (Bai et al. 2020) feeds activation of teacher and student networks to the layers of one another for cross-correction. FSKD (Li et al. 2020b) obtains the student architecture from the teacher itself via pruning, freezes it, and only learns 1x1 convolutions added after each layer. NetGraft (Shen et al. 2021) performs distillation one layer at a time by grafting the student’s layer-to-be-trained to the teacher’s architecture.
Model Inversion for Data-Free KD
If the training dataset is entirely inaccessible, conventional distillation methods can not operate. Data-free methods use the textit Model Inversion (MI) technique to infer data samples that the teacher model had observed during training and use them for distillation. Some early works consider the confidence of teachers’ predictions as supervision for sample generation (Chen et al. 2019). Some others generate samples that maximize the information gain to the student (Micaelli and Storkey 2019). Following these, DeepInversion (Yin et al. 2020) proposed taking advantage of the batch normalization statistics gathered while training the teacher model. CMI (Fang et al. 2021) improved on this method by diversifying sample synthesis with the help of contrastive learning. Fast-Datafree (Fang et al. 2022) proposed a technique to reduce the significant amount of time that data-free distillation takes. PRE-DFKD (Binici et al. 2022) introduced a method to eliminate the trade-off between the large memory footprint and the robustness of the data-free KD process. Recent works have focused on further addressing scalability and efficacy issues (Yu et al. 2023; Liu et al. 2024).
Dataset Condensation
Dataset condensation was introduced by (Zhao, Mopuri, and Bilen 2021) to reduce the training time required for large-scale datasets. It optimizes small batches of synthetic samples to carry almost equal information content as real batches of much larger size. The resulting samples are typically not visually realistic and can better protect data privacy than communicating real samples (Dong, Zhao, and Lyu 2022). To quantitatively assess the privacy benefits, (Zhou, Nezhadarya, and Ba 2022) exercised membership inference attacks (MIA) using condensed samples and showed they yield only around $0.52$ attack AUC, which is almost the same value as random guessing, i.e. $0.5$. Therefore, in cases where samples from a real training set cannot be communicated due to privacy concerns, these condensed samples can be utilized to guide KD’s synthetic data generation process. Moreover, the time cost of dataset condensation has significantly decreased due to recent advancements in the area (Zhao and Bilen 2023; Feng, Vedantam, and Kempe 2023). As such, (Zhou, Nezhadarya, and Ba 2022) can generate a condensed dataset of size 10 samples-per-class from ImageNet-200 in less than an hour with 2 GB memory utilization on a single Nvidia Quadro RTX 6000. This allows condensed versions of datasets at various scales (ranging from MNIST to ImageNet) to be conveniently produced and released to public. (He et al. 2023) further reduces the storage requirement of dataset condensation by compressing multiple condensation processes into a single one.
Condensed Sample-Guided Model Inversion
Model inversion techniques (Yin et al. 2020; Chen et al. 2019; Fang et al. 2021) typically use variants of the following loss function to guide the synthetic data generation.
$\mathcal{L}_{MI}=\mathbb{E}_{x,y_{ps}}\left[y_{ps}log(\hat{y}_{T})\right]-\mathbb{E}_{x}\left[D_{KL}(\hat{y_{T}}||\hat{y_{S}})\right]$
The first term targets to maximize the softmax score ($\hat{y}_{(T)}$) (confidence) that the synthetic samples $\hat{x}$ receive from the teacher $T$ for the pseudo-classes ($y_{ps}\sim u(0,c)$) assigned to them. Generally, $\hat{x}$ is obtained using a generative model parameterized by $\theta_{g}$ ($x\sim p_{\theta_{g}}(x|z)$). The last term encourages the synthesis of samples that provide high information gain to the student. As this objective is strictly guided by the knowledge that teacher embodies, supplementary information cannot be incorporated to provide additional guidance.
Feature alignment mechanism:
To incorporate the supplementary information about the training distribution in the model inversion objective, we add the constraint of aligning the feature distribution of generated synthetic samples with that of the condensed ones. This constraint is enforced through the inclusion of a feature discriminator in our model inversion framework. As illustrated in Figure 2, first, the generator outputs synthetic samples. Then, the teacher model encodes both these synthetic batches and condensed samples into feature representations. Later, discriminator classifies these features as condensed or synthetic samples. This classification results in a feature alignment loss that quantifies the gap between synthetic and condensed feature distributions. This can be viewed as a minimax game in which the feature discriminator competes against the generator for distinguishing real features from synthetic ones. In equilibrium, the generator will be able to provide samples that can yield similar features as the real ones to trick the discriminator. Essentially, the condensed samples serve as prototypes, guiding the generation of new samples that reflect the training distribution.
Figure 2: Overview of our condensed-samples guided model inversion framework.
The final optimization objective for the generator is constructed by combining the feature alignment loss with the objective of any base model inversion method $\mathcal{L}_{MI}$ as shown in Equation 2.
[…]
Where $\mathcal{L}_{FA}=\mathbb{E}_{\hat{x}}\left[1-\log D(\phi_{l}(\hat{x}))\right]$ stands for feature alignment loss. As the dimension of feature vectors is high with respect to the limited availability of condensed samples, our method is prone to over-fitting. Therefore we use a simple discriminator architecture introduced by (Li et al. 2020a) that contains very few parameters. To further address the risk of over-fitting, we perform differentiable data augmentation (Zhao et al. 2020) on both the synthetic images output by the generator and the condensed samples before feeding them to the discriminator.
In deciding on the layer index at which the feature alignment will be employed, we considered the type of image features encoded by different parts of the teacher model. Typically, for image inputs, early layers of neural networks encode structural patterns that are commonly shared across natural images (e.g., edges). In contrast, the image features occurring at the later layers contain semantical information. As our objective is to produce diverse views of objects from the same semantical classes as the condensed samples, we considered features at late layers, specifically the penultimate layer, of the teacher as alignment targets.
Additionally, we posit that simply aligning the cumulative distribution of synthetic features from all classes with the real feature distribution is not ideal. Rather, we provide class-specific alignment using a conditional discriminator (Mirza and Osindero 2014). The contrast between these two alternatives can be seen in the figure given in the appendix. This further changes our discriminator objective to the following.
$\begin{split}\max\mathcal{L}_{D}&=\mathbb{E}_{(x,y)}\left[\log D(\phi_{l}(x),y)\right]\\ &+\mathbb{E}_{(\hat{x},y_{ps})}\left[1-\log D(\phi_{l}(\hat{x}),y_{ps})\right]\end{split}$ (3)
Here, the discriminator not only predicts if a feature is associated with a condensed or synthetic sample but also determines the class it belongs to. To enforce this, we present three different types of inputs to the discriminator. First, we construct “real” inputs by pairing real features with their labels. Later, we use the teacher to assign labels to the synthetic features and obtain “fake” inputs. Lastly, to prevent the discriminator from neglecting the class information, we construct additional “fake” inputs by pairing the same real features with the wrong class labels. Formally, our real ($\mathcal{R}$) and fake ($\mathcal{F}$) sets can be defined as,
=$\begin{split}\mathcal{R}&=\{(x,y)|(x,y)\in\mathcal{X}\}\\ \mathcal{F}&=\{(\hat{x},y_{ps})\}\cup\{(x,c)|(x,y)\in\mathcal{X},c\neq y\}\end{split}$ (4)
Combining condensed and synthetic samples:
After establishing our feature-alignment strategy to improve model inversion with the available data samples, we discuss how we can join condensed and generated synthetic samples for KD. Some alternatives included pre/post-training the student with condensed samples with respect to model inversion. However, as these methods can cause the student to be biased towards one data type, we avoided them. Instead, we extended the condensed dataset by adding the iteratively refined synthetic samples (through model inversion) and trained the student with randomly sampled batches from such union. Our distillation objective involves minimizing the distance between the predictions of the teacher and the student models, which can be summarized as,
$\theta^{*}_{S}:=\arg\min\limits_{\theta_{S}}\mathbb{E}_{\hat{x}}\left[D_{KL}(\hat{y}_{S}||\hat{y}_{T})\right]+\mathbb{E}_{x}\left[D_{KL}(y_{S}||y_{T})\right]$
In Equation 5, $\hat{x}$ and $x$ denote synthetic samples and condensed samples respectively. The exact procedure we follow in generating synthetic samples and distilling the student is summarised in Algorithm 1. First, we initialize our synthetic dataset $\mathcal{X}$ with the available condensed samples. Later at each epoch, we generate a new synthetic batch via our condensed sample-guided model inversion and add it to $\mathcal{X}$. Later, within the same epoch, we randomly draw a data batch from $\mathcal{X}$ and use it to transfer knowledge from the teacher to the student.
Algorithm 1 Knowledge Distillation
[…]
  OUTPUT: trained student $\theta^{*}_{S}$.
  for number of epochs do
     $\hat{x}_{new}\leftarrow invert\_model(T,S,G,D,\mathcal{X})$
     $\mathcal{X}\leftarrow\mathcal{X}\cup\hat{x}_{new}$
     $(x,y)\sim\mathcal{X}$
     $\mathcal{L}_{KD}\leftarrow\sum\limits_{\mathcal{X}}D_{KL}(\hat{y}_{S}||\hat{y}_{T})$
     $optimizer.step(backward(\mathcal{L}_{KD}),\theta_{S})$
  end for
Experimental Evaluation
To assess the effectiveness of our method, we incorporate condensed data guidance to the model inversion procedures of three state-of-the-art data-free KD methods and record the improvement. These methods are Fast, CMI, and PRE-DFKD. Moreover, we also experiment with using limited real samples for feature regularization and observe the advantage against few-shot KD methods. For this comparison, we selected NetGraft and FSKD as baselines. All the results we report on the performance of our baseline methods are either directly taken from the papers or obtained by running the official implementations based on the hyper-parameter configurations shared in the papers or GitHub pages. To standardize the evaluation, we use fixed random seeds borrowed from the official implementations of the baselines.
Datasets
We use three image classification datasets, which are CIFAR-10/100 (Krizhevsky and Hinton 2009), and ImageNet-200 (Deng et al. 2009). We conducted our experiments using condensed samples generated by three different methods, including those provided by (Zhao and Bilen 2021), (Cazenavette et al. 2022) and (Zhao and Bilen 2023). While the results reported throughout our experiments primarily utilize the condensed samples from (Zhao and Bilen 2021), we also include an ablation study to compare the effectiveness of each condensation method.
Implementation details
We used the same generator architectures as shared in the official implementations of the data-free methods that we couple our method with (Fast, CMI, and PRE-DKD). Further details on generator and discriminator architectures as well as how we couple our method with individual MI methods can be found in the appendix.
Condensed-data guided sample synthesis
Tables 1 and 4 show the student accuracies upon coupling our approach with Fast, CMI, and PRE-DFKD. The results achieved by this coupling are indicated by the asterisk symbol with the annotation “*(ours w/ CS)”. To eliminate the advantage of having access to condensed samples, we also configure baselines where we combine these samples with the synthetic datasets generated by model inversion methods. These baselines are denoted with “+ CS” notations. Further, the row “Train w/ full real data” represents the accuracy of full-scale training of students on the target dataset and constitutes the upper bound. “Train w/ cond. samples” reflects the accuracy achieved by only using condensed samples for student training, which is the lower bound. In all experiments, we assume 50 spc (500 samples) and 10 spc (1000 samples) are available for CIFAR-10 and CIFAR-100, respectively. For ImageNet-200, we consider the availability of 10 spc (2000 samples). These represent 1%, 2%, and 2% of the total samples in their respective datasets.
Dataset CIFAR-10 CIFAR-100 ImageNet-200
Teacher ResNet-34 ResNet-34 WRN-40-2 WRN-40-2 ResNet-34 ResNet-34 WRN-40-2 WRN-40-2 ResNet-34 ResNet-34
Student ResNet-18 MBNet-v2 WRN-40-2 MBNet-v2 ResNet-18 MBNet-v2 WRN-40-2 MBNet-v2 ResNet-18 MBNet-v2
Teacher acc. 95.70 95.70 94.87 94.87 78.05 78.05 75.83 75.83 71.20 71.20
Training student (S) with labeled data
Train w/ full real data
95.20 93.79 94.87 93.79 77.10 72.80 75.83 72.80 64.90 55.06
Train w/ cond. samples (CS)
34.66 30.13 38.92 30.13 16.54 10.69 15.55 10.69 3.50 5.60
Distilling student (S) with synthetic data
Fast 92.62 86.12 92.82 85.06 69.76 54.62 65.05 48.21 42.99 35.31
Fast + CS 92.72 86.37 92.84 85.69 69.96 56.57 65.51 49.42 45.43 40.68
Fast* (ours w/ CS) 94.33 88.05 94.64 89.24 72.09 63.29 70.96 60.86 48.14 43.08
CMI 94.84 87.5 92.83 86.53 77.04 61.9 68.96 59.04 44.11 35.55
CMI + CS 94.89 88.06 92.94 86.77 77.04 62.54 69.10 59.62 47.07 40.67
CMI* (ours w/ CS) 94.97 89.63 94.21 90.21 77.07 70.21 72.42 68.05 48.98 45.83
Table 1: Student accuracies (%) obtained by enhancing Fast and CMI with condensed-sample guidance. Condensed datasets with 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200 were used.
((a)) Synthetic feature distribution by data-free model inversion
((b)) Condensed data-guided synthetic feature distribution (ours)
Figure 3: 2D visualisation of feature vectors. Faded and bold markers denote feature space projections of real samples from the CIFAR-10 dataset and synthetic samples, respectively. Condensed data-guided synthetic samples exhibit better alignment with the real data distribution.
First, we note that the performance of Fast method notably diminishes for pairs with low structural similarity (e.g. ResNet-34 (He et al. 2016) & MobileNet-v2 (Howard et al. 2017)). The improvement was especially significant for WRN-40-2 (Zagoruyko and Komodakis 2016b) & MobileNet-v2 pairs reaching up to $11.44\%$ on CIFAR-100. CMI also has a considerable performance gap with respect to the upper bound for heterogeneous model pairs. Our method again achieves consistent advantage, with substantial accuracy improvements reaching up to $8.43\%$ (WRN-40-2 & MobileNet-v2). Since PRE-DFKD achieves almost the same student accuracy as the upper limit for homogeneous pairs (e.g. ResNet-34 & ResNet-18), we only experiment with heterogeneous ones (e.g. ResNet-34 & MobileNet-v2), where there is still room for improvement. The results in Table 4 shows that our method effectively improves acuracy also for this baseline method.
In all experiments, the simple inclusion of condensed samples in the distillation set (“+ CS”) neither mitigated this issue nor caused any substantial performance improvement in most cases. On the other hand, our condensed sample-guided model inversion (“*”) consistently increased student accuracy across different datasets and teacher-student pairs, which ensures that the benefit of our approach is not simply due to exposing the student to condensed samples during training.
Visual results
We examine the impact of our method on the visual quality of the generated samples. Figure 4 contains synthetic CIFAR100 images obtained by CMI and CMI*. We note that CMI* samples are significantly more realistic and exhibit common class-distinctive patterns across images from the same categories, which is not observed in CMI. This strengthens the claim that feature alignment effectively conditions the synthetic data to contain realistic semantics that is consistent among samples from the same classes.
Figure 4: First two rows contain synthetic CIFAR-100 samples obtained with and w/o condensed data-guided model inversion. The last two show condensed and real samples.
Figure 5: Students distilled by vanilla and condensed sample-guided model inversion for different condensed dataset set sizes. Condensed datasets with 1, 10, and 50 spc from CIFAR10, and 1 and 10 spc from CIFAR100, were used.
Additionally, this conditioning does not compromise the diversity of the synthetic set, as demonstrated by the varied object views and scales in CMI*. We also note that neither the condensed samples nor the condensed sample-guided generated synthetic samples exist in the real dataset. Thus, although the generated synthetic samples are visually realistic, they do not reveal any individual training samples. Moreover, to visually observe the effect of our method on the distribution of generated synthetic samples, we projected the high-dimensional feature space on a 2D plane using the t-SNE algorithm (Van der Maaten and Hinton 2008). On the resulting plane, we compare the feature distributions observed for CMI and CMI*. We use the WRN-40-2 teacher trained on the CIFAR-10 dataset as the feature extractor. The visualizations are displayed in Figure 3 where faded data points mark the features of real samples, and the bold ones represent the synthetic ones. Each sample is marked with the color associated with its class label. We observe that synthetic samples generated by CMI are mostly clustered together and only partially correspond with real samples. Additionally, some feature clusters have almost no correspondence with real features (circled in red), meaning that the associated synthetic samples might not be contributing to the knowledge transfer. In contrast, the synthetic feature distribution formed during CMI* shows better alignment with real data distribution.
How does student accuracy scale with the condensed data size?
After establishing that our proposed method can boost performance using condensed samples, we analyze how the scale of the available data affects the improvement. For this, we consider condensed datasets of 3 different sizes (1 spc, 10 spc, 50 spc) for CIFAR-10 and 2 different sizes (1 spc, 10 spc) for CIFAR-100. These amounts correspond to $0.02\%$, $0.2\%$ and $1\%$ of the samples contained in CIFAR-10, and $0.2\%$ and $2\%$ for CIFAR-100. The plots in Figure 5 show that Fast and CMI baselines do not benefit from the mere inclusion of condensed samples in their synthetic distillation sets, irrespective of the number of samples available. However, when they are equipped with our feature alignment module, the student accuracies scale up with increasing data availability.
Does the condensation method affect the quality of model inversion?
We study the impact of the dataset condensation method used to produce the condensed samples on the effectiveness of our approach and display the results in Table 2. DSA, DM and MTT refer to (Zhao and Bilen 2021), (Zhao and Bilen 2023), and (Cazenavette et al. 2022) respectively. The results indicate that while the choice of condensation method can impact the final student accuracy, our method consistently improves performance across all types of condensed samples tested. This suggests that its effectiveness is not dependent on any single condensation approach.
Dataset CIFAR-10
Teacher ResNet-34 WRN-40-2
Student MobileNet-v2 MobileNet-v2
Ablation Study on Different Dataset Condensation Methods
KD Method CS Type DSA DM MTT DSA DM MTT
Fast + CS 86.37 86.54 86.78 85.69 80.01 80.38
Fast* (ours w/ CS) 88.05 87.82 87.11 89.24 90.30 90.38
Table 2: Impact of different types of condensed samples (CS) on student accuracy (%) for different dataset condensation strategies. Condensed dataset with 50 spc from CIFAR10 was used.
Ablation study
To demonstrate the individual contributions of the constituents of method to the final performance, we provide an ablation study in Table 2.
Method Student acc.
CMI 92.83
+ CS 92.94
+ CS-guided MI 93.98
+ class-specific alignment (CMI*) 94.21
Table 3: MobileNet-v2 student accuracy on CIFAR10 for various configurations of the components used in our method. Condensed dataset with 50 spc from CIFAR10 was used.
Dataset CIFAR-10 CIFAR-100 ImageNet-200
Teacher ResNet-34 WRN-40-2 ResNet-34 WRN-40-2 ResNet-34 ResNet-34
Student MobileNet-v2 MobileNet-v2 MobileNet-v2 MobileNet-v2 ResNet-18 MobileNet-v2
Teacher acc. 95.70 94.87 78.05 75.83 71.20 71.20
Accuracy of student (S) trained with labeled data
Train w/ full real data 93.79 93.79 72.80 72.80 64.90 55.06
Train w/ cond. samples (CS) 27.24 27.24 10.69 10.69 3.50 5.60
Train w/ few real samples (RS) 27.55 27.55 4.66 4.66 1.42 1.30
Accuracy of student (S) trained with synthetic data
PRE-DFKD 83.12 83.12 66.56 61.83 54.20 47.26
PRE-DFKD* (ours w/ CS) 86.89 88.34 67.39 63.41 55.22 49.73
PRE-DFKD* (ours w/ RS) 86.77 87.47 71.63 64.37 54.95 49.68
Table 4: Student accuracies (%) obtained by enhancing PRE-DFKD with condensed and real-sample guidance (CS and RS). Condensed and real sample sets containing 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200 were used.
Dataset CIFAR-10 CIFAR100 Teacher VGG-11 VGG-11 VGG-11 VGG-11 Student VGG-11 (50% pruned) VGG-11 (75% pruned) VGG-11 (25% pruned) VGG-11 (50% pruned) Teacher acc. 92.25 92.25 71.23 71.23 Few-shot distillation accuracy FSKD 78.69 36.00 54.72 24.73 PRE-DFKD* 83.26 68.48 63.81 61.21 ((a)) Student accuracy comparison with FSKD baseline. Real sample sets containing 50 spc from CIFAR10, and 10 spc from CIFAR100 were used. Dataset CIFAR-10 CIFAR100 Teacher VGG-16 VGG-16 Student ResNet-18 ResNet-18 Teacher acc. 94.16 74.00 Few-shot distillation accuracy NetGraft 73.69 55.51 PRE-DFKD* 88.38 68.54 ((b)) Comparison with few-shot NetGraft baseline. 1 spc from both CIFAR10 / 100 were used.
Table 5: Comparison with different few-shot baselines using few real samples.
This study involves student accuracies upon enhancing a base model inversion method (CMI) in 3 different settings. (1) utilizing condensed samples only by appending them to the distillation set (“+ CS”), (2) using them also to regularize the feature distribution of synthetic data (“+ CS-guided MI”) (3) the final version of our condensed sample-guided model inversion (CMI*). As seen in Table 3 “+ CS-guided MI” achieves advantage over simply including condensed samples in the distillation set (“+ CS”). Later, we include the final component, i.e., class-specific alignment, and record student accuracy (“+ class-specific alignment”). Although this addition caused only a marginal increase in performance, we still retain it as it improves results without introducing any notable computational burden.
Real-data guided sample synthesis
Our method is also inherently capable of utilizing limited real samples from the target dataset when provided. To show this, we repeated the experiments in Table 4 by replacing the condensed samples with real ones. We assume the availability of the same amount of samples randomly drawn from the target datasets as we used in experiments with condensed data (50 spc for CIFAR10, 10 spc for CIFAR100 and 10 spc for ImageNet-200). The results are displayed in Table 4 as PRE-DFKD* (ours w/ RS). Similar to condensed sample experiments, the improvements observed for heterogeneous pairs are again greater than the homogeneous ones, i.e. ResNet-34 & ResNet-18. The overall improvement yielded by our method upon utilizing few real samples is comparable that using condensed ones. This highlights the effectiveness of our approach in few-shot KD scenarios.
Comparison with few-shot methods
We benchmarked the effectiveness of our method in few-shot KD against FSKD and Netgraft baselines using the same experiment setups reported in their papers. To ensure a fair evaluation of our work against these few-shot baselines, we selected identical teacher-student pairs to those in the original papers. We again use the same amount of samples from both datasets as in Table 4. In our comparison with FSKD, we use VGG11 teachers and student models that are channel-pruned versions of the teacher at different rates. As for the comparison with Netgraft, we use VGG16 teachers and ResNet-18 students, with 1 spc from both datasets. From Table 5, it can be observed that FSKD is only effective for low pruning rates while performing poorly for higher rates. Similarly, the accuracies achieved by NetGraft were also much lower than the upper-bound as seen in Table 5. This is expected as both baselines are restricted to the limited information made available by the few-shot sample set. As our method has a generative component, it does not have such restriction and therefore outperforms both few-shot baselines by a large margin.
Conclusion
In conclusion, our study addresses the challenge of enhancing data-free knowledge distillation by incorporating supplementary information available from the target dataset—something traditional model inversion methods fail to leverage. Recognizing the potential of condensed samples to serve as this supplementary information, we developed a method that refines the synthetic data generation process within the model inversion framework. This integration helps align the synthetic data more closely with the real data distribution, effectively reducing the domain gap that often undermines KD performance. Our experiments demonstrate the effectiveness of this approach, showing accuracy improvements of up to $11.4\%$ across various datasets and model configurations. The method is also applicable in environments where only a few real samples are accessible, such as in few-shot learning scenarios.
Acknowledgement
This work is supported by the National Research Foundation, Singapore under its Competitive Research Programme Award NRF-CRP23-2019-0003 and Singapore Ministry of Education Academic Research Fund T1 251RES1905. We thank Cihan Acar and Liu Siying for helpful comments on the manuscript.
Appendix A: Computing Infrastructure
The specifications of the system used to run computational experiments are:
  • Ubuntu 22.04.4 LTS
  • NVIDIA A100 80GB GPUs
  • Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz
Appendix B: Implementation Details
Generator architectures
The Fast and CMI baselines share the same generator architecture, which is detailed in Table 6. The PRE-DFKD generator differs slightly, with the inclusion of spectral normalization following the convolution operations, as outlined in Table 7. In addition, the final activation function is hyperbolic tangent (Tanh), as opposed to sigmoid. Both tables utilize ”$nz$” to represent the dimension of the random vector that is drawn from a standard normal distribution. In PRE-DFKD, the value of ”$nz$” is set to 1000, while in Fast and CMI, it is set to 256. Lastly, the described architectures are cofigured to produce 32x32 RGB images (for CIFAR-10 and CIFAR-100). The activation shapes vary for datasets with samples of higher resolution (e.g. ImageNet-200).
Layer # Operation in shape out shape
1 Linear nz 8x8x128
2 BN 8x8x128 8x8x128
3 2xUpsample 8x8x128 16x16x128
4 Conv-BN-ReLU 16x16x128 16x16x128
5 2xUpsample 16x16x128 32x32x128
6 Conv-BN-ReLU 32x32x128 32x32x64
7 Conv 32x32x64 32x32x3
8 Sigmoid 32x32x3 32x32x3
Table 6: Generator architecture used in Fast and CMI experiments
Layer # Operation in shape out shape
1 Linear nz 8x8x128
2 BN 8x8x128 8x8x128
3 2xUpsample 8x8x128 16x16x128
4 Conv-SN-BN-ReLU 16x16x128 16x16x128
5 2xUpsample 16x16x128 32x32x128
6 Conv-SN-BN-ReLU 32x32x128 32x32x64
7 Conv-SN 32x32x64 32x32x3
8 Tanh 32x32x3 32x32x3
9 BN 32x32x3 32x32x3
Table 7: Generator architecture used in PRE-DFKD experiments. SN stands for spectral normalization.
Discriminator architecture & training details
For all experiments, we utilize the same discriminator whose architecture is described in Table 3. “nz” has the same meaning as in Tables 6 and 7. “nc” corresponds to the number of classes, for instance “nc=100” for CIFAR-100 dataset. To update the discriminator we use Adam optimizer with learning rate=$0.002$ and beta values=$[0.5,0.999]$. While propagating the discriminator loss to the generator, we scale it with the coefficient of $0.5$, thus the generator loss $L_{G}$ in Equation 2 becomes:
${L}_{G}=\mathcal{L}_{MI}+0.5\times\mathcal{L}_{FA}$ (6)
Layer # Operation in shape out shape
1 Linear nz+nc 1024
2 ReLU 1024 1024
3 Dropout (0.5) 1024 1024
4 Linear 1024 1
Table 8: Architectural details of the feature discriminator.
Coupling with model inversion methods
Fast and CMI use only a single generator, so we simply attach our discriminator as shown in Figure 1(a) in the main paper. In addition, we notice that both methods reset the generator periodically by randomly re-initializing its weights. To prevent any instability in our adversarial feature regularization setup, we also reset the discriminator with the same frequency. We employ the Fast-2 configuration on CIFAR-10 and CIFAR-100 datasets, which update the generator two times for every update that the student receives. Meanwhile, for ImageNet-200, we adhere to the Fast-50 configuration, which is consistent with the original paper’s approach for complex datasets. As the PRE-DFKD framework contains 2 generators, namely “novel” and “memory” generators, attaching our method requires further consideration. While the “novel” generator is responsible for producing new/informative samples in the framework, the memory generator memorizes the cumulative output distribution of the novel generator. As conditioning the novel generator also implicitly affects the memory generator, we only pair the discriminator with it. We run CMI, Fast, and PRE-DFKD for 200, 220, and 220 epochs, respectively. All results reflect the maximum student accuracy observed during distillation.
Moreover, PRE-DFKD does not store generated samples and instead distills the student on-the-fly with newly generated samples at each epoch. This prevents us from directly applying Algorithm 1 to combine condensed and synthetic samples, and the other alternatives discussed in Section 3 impair student accuracy. For this reason, we only use the condensed samples to guide synthetic data generation and exclude them while training the student. Even without utilizing condensed samples in a direct way in distillation, we observe accuracy improvement on both datasets.
Appendix C: Condensed Datasets
The basic dataset condensation process uses a reference model and optimizes condensed samples to match the gradients that larger real batches produce for that model. In our experiments, we used condensed samples retrieved from the official GitHub pages of (Zhao, Mopuri, and Bilen 2021) and (Cazenavette et al. 2022). Both works, used the same 3-layer CNN architecture as the reference model to generate these datasets. For CIFAR-10, we used samples generated by both (Zhao and Bilen 2021) and (Zhao and Bilen 2023). The samples from (Zhao and Bilen 2021) are available in three sizes: 1, 10, and 50 samples per class (spc), while (Zhao and Bilen 2023) provides samples only at 50 spc. For CIFAR-100, only the 1 and 10 spc settings from (Zhao and Bilen 2021) are available. For ImageNet-200, we utilized the condensed samples from (Cazenavette et al. 2022), which provides 10 spc. It is worth noting that none of the teacher models in our experiments were used as reference models during the generation of these samples. This demonstrates that our method can effectively utilize condensed datasets, regardless of the reference model used in their optimization process.
Appendix D: Time complexity of model inversion
To show that our method does not sacrifice significant speed while enhancing the performance, we share the model inversion times in table 9.
Method Time (hrs)
Fast 3
Fast* 4
CMI 11
CMI* 13.5
PRE-DFKD 11.6
PRE-DFKD* 13.5
Table 9: Times consumed by different model inversion methods.
Appendix E: Class-specific Feature Alignment
When the discriminator is configured to only differentiate between the generated synthetic samples and the condensed ones, the generator can achieve its optimization objective by merely aligning the two distributions, as demonstrated in Figure 6(a). However, for optimal results, it is preferable to align the synthetic samples with the condensed ones based on their respective classes, as shown in Figure 6(b). For instance, a synthetic car image should exhibit similar semantic features to a condensed car image rather than a condensed sunflower image.
((a)) generic alignment
((b)) class-specific alignment
Figure 6: Illustration of feature spaces observed after different feature alignment strategies. Condensed (C) and synthetic (S) features are denoted by red and blue colors respectively. Each shape indicates a unique class ID (e.g. c1).

### v2

@@
-Condensed Sample-Guided Model Inversion for Knowledge Distillation
+Condensed Data Expansion Using Model Inversion for Knowledge Distillation
 Abstract
-Knowledge distillation (KD) is a key element in neural network compression that allows knowledge transfer from a pre-trained teacher model to a more compact student model.
-KD relies on access to the training dataset, which may not always be fully available due to privacy concerns or logistical issues related to the size of the data.
-To address this, ”data-free” KD methods use synthetic data, generated through model inversion, to mimic the target data distribution.
-However, conventional model inversion methods are not designed to utilize supplementary information from the target dataset, and thus, cannot leverage it to improve performance, even when it is available.
-In this paper, we consider condensed samples, as a form of supplementary information, and introduce a method for using them to better approximate the target data distribution, thereby enhancing the KD performance.
-Our approach is versatile, evidenced by improvements of up to $11.4\%$ in KD accuracy across various datasets and model inversion-based methods.
+Condensed datasets offer a compact representation of larger datasets, but training models directly on them or using them to enhance model performance through knowledge distillation (KD) can result in suboptimal outcomes due to limited information.
+To address this, we propose a method that expands condensed datasets using model inversion, a technique for generating synthetic data based on the impressions of a pre-trained model on its training data.
+This approach is particularly well-suited for KD scenarios, as the teacher model is already pre-trained and retains knowledge of the original training data.
+By creating synthetic data that complements the condensed samples, we enrich the training set and better approximate the underlying data distribution, leading to improvements in student model accuracy during knowledge distillation.
+Our method demonstrates significant gains in KD accuracy compared to using condensed datasets alone and outperforms standard model inversion-based KD methods by up to $11.4\%$ across various datasets and model architectures.
 Importantly, it remains effective even when using as few as one condensed sample per class, and can also enhance performance in few-shot scenarios where only limited real data samples are available.
 Introduction
-Knowledge distillation (KD) (Hinton, Vinyals, and Dean 2015) is an effective technique for reducing the resource demand of large pre-trained models.
-It functions by transferring the task knowledge of a large pre-trained model called the “teacher” to a more compact “student” architecture.
-One restrictive aspect of KD is its dependence on the training dataset of the target task.
-Such data might not be available in full scale, making distillation ineffective or even infeasible.
-For example, the training data could be privacy sensitive or large in memory size, aggravating its relocation.
-Addressing this issue, several works examine methods to perform knowledge distillation even when no sample from the training set is available at the compression stage.
-These are called data-free distillation techniques and are mainly based on generating synthetic samples to substitute the real training set (Lopes, Fenu, and Starner 2017; Yu et al.
-2023; Liu et al.
-2024).
-To generate informative samples that are aligned with the real data distribution, the information encapsulated in the teacher model is leveraged, which is called model inversion.
+Condensed datasets (zhao2021dataset) have emerged as a promising approach for compactly representing large datasets, enabling efficient model training with reduced memory and computational costs.
+These datasets consist of synthetic samples optimized to capture the information content of much larger datasets.
+They provide certain privacy benefits, as studied in (dong2022privacy) and can be produced with modest memory and time resources through recent methods (zhou2022dataset; zhao2023dataset; feng2023embarrassingly).
+These qualities render condensed samples suitable for scenarios in which the privacy considerations prohibit exposure of individual training samples or the large memory size aggravating their relocation.
+However, the utility of condensed datasets at small scale can be limited in various learning paradigms (yu2023dataset) such as standard supervised learning or knowledge distillation (KD) (hinton2015distilling).
+This limited utility stems from the reduced information captured in these compact representations, hindering the ability of models to effectively learn from them.
+In KD, this limitation is particularly relevant as a student model learns from the guidance of a pre-trained teacher model, whose ability to transfer knowledge is itself constrained by the limited information present in the condensed dataset.
 Figure 1: Illustration of our motivation for using condensed samples as prototypes for synthetic data.
-In certain situations, some information from the target dataset can be communicated.
-However, as model inversion technique is primarily designed to facilitate knowledge distillation when no training data is available, even when such information is supplied, existing methods cannot leverage it for enhancing KD performance.
-In this work, we explore the potential of condensed samples (Zhao, Mopuri, and Bilen 2021) from target data as a form in which supplementary information can be available, and introduce a method to refine the synthetic data generation process in model inversion whenever they are available.
-Our method uses condensed sampled as prototypes to propagate new samples from the training distribution as illustrated in Figure 1.
-Condensed samples are synthetic samples that summarize large batches from target dataset.
-They provide certain privacy benefits, as studied in (Dong, Zhao, and Lyu 2022) and can be produced with modest memory and time resources through recent methods (Zhou, Nezhadarya, and Ba 2022; Zhao and Bilen 2023; Feng, Vedantam, and Kempe 2023).
-These qualities render condensed samples suitable for scenarios in which the privacy considerations prohibit exposure of individual training samples.
-A trivial approach is simply combining the condensed samples with those obtained by model inversion to distill the student.
-However, preliminary experiments suggest that this does not improve accuracy, which is likely caused by the domain gap between synthetic and condensed samples.
-With this observation, we propose using them to refine the quality of synthetic samples generated by model inversion instead.
-Our approach mainly targets improving the assumptions made by model inversion methods about the target data distribution.
-One such assumption is that real samples will be classified by the pre-trained teacher with high confidence, receiving a one-hot-like prediction vector.
-However, in many cases, models can give less confident predictions to real samples, yet still classify them correctly.
-Thus, guiding the generation process based on such assumptions on the confidence of teacher’s predictions can produce synthetic samples that do not follow the real training distribution.
-Our method utilizes a small set of condensed samples to query the teacher model and extract more realistic impressions related to the target dataset.
-Specifically, these samples are fed to the teacher model for estimating the per-class feature distributions of the target data.
+In this work, we address the limited utility of condensed datasets, particularly in KD, by expanding them using model inversion (MI) (lopes2017data; yu2023data; liu2024small).
+Our goal is to enhance the limited information captured by these compact representations and better approximate the underlying training data.
+MI is a technique that leverages a pre-trained model as a discriminator to generate synthetic samples resembling real ones.
+This requirement of a pre-trained model is naturally satisfied in KD, as it inherently involves a pre-trained teacher model.
+MI operates by training a generative model to produce data points that follow the learned distribution of the teacher’s representations (liu2021data).
+Since the true data distribution is unknown without access to the original dataset, the generative model is optimized based on certain inductive biases (zhao2018bias; goyal2022inductive) about the training data.
+One such bias is the assumption that the teacher classifies real samples with high confidence, resulting in near one-hot prediction vectors (chen2019data; yin2020dreaming).
+A trivial approach for expanding the condensed data through MI is simply combining it with the synthetic samples obtained by MI.
+However, our preliminary experiments suggest that this does not improve accuracy, which is likely caused by the domain gap between the two sample sets (hennicke2024mind; bai2024bridging).
+With this observation, in this paper we propose using the condensed samples as prototypes that represent the real data distribution and changing the MI process to generate samples that are more aligned with the characteristics of the condensed data.
+By prioritizing the generation of synthetic samples that resemble these prototypes, our method can bridge the domain gap and enhance the performance of the student model trained on the expanded data.
+Another reason contributing to such performance improvement is that the inductive bias used by MI to model the real data distribution is improved with the knowledge of the prototypes.
+Specifically, our method utilizes a small set of condensed samples to query the teacher model and extract more realistic impressions related to the target dataset.
+These samples are fed to the teacher model for estimating the per-class feature distributions of the target data.
 Then, the model inversion process is conditioned to produce synthetic samples following a similar feature distribution as the condensed ones.
-This is achieved by configuring a feature discriminator that competes against the sample generator to distinguish condensed samples from the ones generated by model inversion.
+This is achieved by configuring a feature discriminator (li2020adversarial) that competes against the sample generator to distinguish condensed samples from the ones generated by model inversion.
 By doing so, the generator is forced to produce synthetic samples with semantic similarity to the condensed ones, thus reducing the risk of a domain gap and improving accuracy.
@@
 In addition, remarkably, even using as few as one condensed sample per class results in a noticeable accuracy improvement.
-Moreover, our method is also applicable to scenarios where a limited amount of real samples from the training set are accessible, such as in few-shot learning scenarios.
+Moreover, our method is also applicable to scenarios where a limited amount of real samples from the training set are accessible, such as in few-shot learning scenarios (song2023comprehensive; sauer2022knowledge).
 Experimental results exhibit the advantage of using synthetic data samples generated with guidance from the real samples against pure few-shot KD methods.
-In other words, if condensed samples are available, our method can use them to enhance the performance of model inversion-based KD (data-free KD).
-Alternatively, it can also enhance KD performance if only a few samples from the real dataset are available.
-Otherwise, it simply reduces to any data-free KD method that it is coupled with.
-Lasty, our method does not introduce a significant time overhead, which is limited to the training cost of our lightweight discriminator.
-Our contributions in this work can be summarized as follows.
-  • A novel approach to take advantage of the supplementary information that can be available from target data, which can be plugged into any model inversion method.
-  • Investigating the applicability of our approach where either condensed samples or a few real samples from target data can be provided.
 Related Work
+Dataset Condensation
+Dataset condensation was introduced by (zhao2021dataset) to reduce the training time required for large-scale datasets.
+It optimizes small batches of synthetic samples to carry almost equal information content as real batches of much larger size.
+The resulting samples are typically not visually realistic and can better protect data privacy than communicating real samples (dong2022privacy).
+To quantitatively assess the privacy benefits, (zhou2022dataset) exercised membership inference attacks (MIA) using condensed samples and showed they yield only around $0.52$ attack AUC, which is almost the same value as random guessing, i.e.
+$0.5$.
+Therefore, in cases where samples from a real training set cannot be communicated due to privacy concerns, these condensed samples can be utilized to guide KD’s synthetic data generation process.
+Moreover, the time cost of dataset condensation has significantly decreased due to recent advancements in the area (zhao2023dataset; feng2023embarrassingly).
+As such, (zhou2022dataset) can generate a condensed dataset of size 10 samples-per-class from ImageNet-200 in less than an hour with 2 GB memory utilization on a single Nvidia Quadro RTX 6000.
+This allows condensed versions of datasets at various scales (ranging from MNIST to ImageNet) to be conveniently produced and released to public.
+(hemultisize) further reduces the storage requirement of dataset condensation by compressing multiple condensation processes into a single one.
+Figure 2: (a) Overview of our condensed-samples guided GMI (generative model inversion) framework.
+The discriminator is optimized to distinguish real and fake features, while the generator tries to prevent it from doing so by aligning them.
+Generator, discriminator, and student models are trained in alternate steps.
+(b) Illustration of our motivation for using condensed samples as templates for synthetic data.
 Knowledge Distillation (KD)
-KD (Hinton, Vinyals, and Dean 2015) trains a compact “student” neural network model to approximate the decision space of a more complex one called the “teacher”.
+KD (hinton2015distilling) trains a compact “student” neural network model to approximate the decision space of a more complex one called the “teacher”.
 The inclusion of the soft guidance supplied by the teacher enriches the limited information the student receives from the one-hot encoded class labels.
 As a result, the student can achieve better performance than supervised training alone can provide.
-While most commonly, the logit scores or the softmax probabilities of the teacher are considered for regularization (Romero et al.
-2014), activation maps or attention scores can also be used (Zagoruyko and Komodakis 2016a).
+While most commonly, the logit scores or the softmax probabilities of the teacher are considered for regularization (romero2014fitnets), activation maps or attention scores can also be used (zagoruyko2016paying).
+Model Inversion
+If the training dataset is entirely inaccessible, conventional KD methods can not operate.
+The textit Model Inversion (MI) technique is developed to infer data samples that the teacher model had observed during training and use them for distillation.
+Some early works consider the confidence of teachers’ predictions as supervision for sample generation (chen2019data).
+Some others generate samples that maximize the information gain to the student (micaelli2019zero).
+Following these, DeepInversion (yin2020dreaming) proposed taking advantage of the batch normalization statistics gathered while training the teacher model.
+CMI (fang2021contrastive) improved on this method by diversifying sample synthesis with the help of contrastive learning.
+Fast-Datafree (fang2022up) proposed a technique to reduce the significant amount of time that model inversion takes.
+PRE-DFKD (binici2022robust) introduced a method to eliminate the trade-off between the large memory footprint and the robustness of the process.
+Recent works have focused on further addressing scalability and efficacy issues (yu2023data; liu2024small).
 Few-Shot KD
@@
 To avoid over-fitting, available methods typically reduce the number of parameters required to train the student network.
-(Bai et al.
-2020) feeds activation of teacher and student networks to the layers of one another for cross-correction.
-FSKD (Li et al.
-2020b) obtains the student architecture from the teacher itself via pruning, freezes it, and only learns 1x1 convolutions added after each layer.
-NetGraft (Shen et al.
-2021) performs distillation one layer at a time by grafting the student’s layer-to-be-trained to the teacher’s architecture.
-Model Inversion for Data-Free KD
-If the training dataset is entirely inaccessible, conventional distillation methods can not operate.
-Data-free methods use the textit Model Inversion (MI) technique to infer data samples that the teacher model had observed during training and use them for distillation.
-Some early works consider the confidence of teachers’ predictions as supervision for sample generation (Chen et al.
-2019).
-Some others generate samples that maximize the information gain to the student (Micaelli and Storkey 2019).
-Following these, DeepInversion (Yin et al.
-2020) proposed taking advantage of the batch normalization statistics gathered while training the teacher model.
-CMI (Fang et al.
-2021) improved on this method by diversifying sample synthesis with the help of contrastive learning.
-Fast-Datafree (Fang et al.
-2022) proposed a technique to reduce the significant amount of time that data-free distillation takes.
-PRE-DFKD (Binici et al.
-2022) introduced a method to eliminate the trade-off between the large memory footprint and the robustness of the data-free KD process.
-Recent works have focused on further addressing scalability and efficacy issues (Yu et al.
-2023; Liu et al.
-2024).
-Dataset Condensation
-Dataset condensation was introduced by (Zhao, Mopuri, and Bilen 2021) to reduce the training time required for large-scale datasets.
-It optimizes small batches of synthetic samples to carry almost equal information content as real batches of much larger size.
-The resulting samples are typically not visually realistic and can better protect data privacy than communicating real samples (Dong, Zhao, and Lyu 2022).
-To quantitatively assess the privacy benefits, (Zhou, Nezhadarya, and Ba 2022) exercised membership inference attacks (MIA) using condensed samples and showed they yield only around $0.52$ attack AUC, which is almost the same value as random guessing, i.e.
-$0.5$.
-Therefore, in cases where samples from a real training set cannot be communicated due to privacy concerns, these condensed samples can be utilized to guide KD’s synthetic data generation process.
-Moreover, the time cost of dataset condensation has significantly decreased due to recent advancements in the area (Zhao and Bilen 2023; Feng, Vedantam, and Kempe 2023).
-As such, (Zhou, Nezhadarya, and Ba 2022) can generate a condensed dataset of size 10 samples-per-class from ImageNet-200 in less than an hour with 2 GB memory utilization on a single Nvidia Quadro RTX 6000.
-This allows condensed versions of datasets at various scales (ranging from MNIST to ImageNet) to be conveniently produced and released to public.
-(He et al.
-2023) further reduces the storage requirement of dataset condensation by compressing multiple condensation processes into a single one.
-Condensed Sample-Guided Model Inversion
-Model inversion techniques (Yin et al. 2020; Chen et al. 2019; Fang et al. 2021) typically use variants of the following loss function to guide the synthetic data generation.
-$\mathcal{L}_{MI}=\mathbb{E}_{x,y_{ps}}\left[y_{ps}log(\hat{y}_{T})\right]-\mathbb{E}_{x}\left[D_{KL}(\hat{y_{T}}||\hat{y_{S}})\right]$
+(bai2020few) feeds activation of teacher and student networks to the layers of one another for cross-correction.
+FSKD (li2020few) obtains the student architecture from the teacher itself via pruning, freezes it, and only learns 1x1 convolutions added after each layer.
+NetGraft (shen2021progressive) performs distill
+Condensed Data Expansion Using MI
+Model inversion techniques (yin2020dreaming; chen2019data; fang2021contrastive) typically use variants of the following loss function to guide the synthetic data generation.
+[…]
 The first term targets to maximize the softmax score ($\hat{y}_{(T)}$) (confidence) that the synthetic samples $\hat{x}$ receive from the teacher $T$ for the pseudo-classes ($y_{ps}\sim u(0,c)$) assigned to them.
@@
 The last term encourages the synthesis of samples that provide high information gain to the student.
-As this objective is strictly guided by the knowledge that teacher embodies, supplementary information cannot be incorporated to provide additional guidance.
+As this objective is strictly guided by the knowledge that teacher embodies, prototype information cannot be incorporated to provide additional guidance.
 Feature alignment mechanism:
-To incorporate the supplementary information about the training distribution in the model inversion objective, we add the constraint of aligning the feature distribution of generated synthetic samples with that of the condensed ones.
+To incorporate the condensed data prototypes form the training distribution in the model inversion objective, we add the constraint of aligning the feature distribution of generated synthetic samples with that of the condensed ones.
 This constraint is enforced through the inclusion of a feature discriminator in our model inversion framework.
@@
 Essentially, the condensed samples serve as prototypes, guiding the generation of new samples that reflect the training distribution.
-Figure 2: Overview of our condensed-samples guided model inversion framework.
 The final optimization objective for the generator is constructed by combining the feature alignment loss with the objective of any base model inversion method $\mathcal{L}_{MI}$ as shown in Equation 2.
 […]
-Where $\mathcal{L}_{FA}=\mathbb{E}_{\hat{x}}\left[1-\log D(\phi_{l}(\hat{x}))\right]$ stands for feature alignment loss.
-As the dimension of feature vectors is high with respect to the limited availability of condensed samples, our method is prone to over-fitting.
-Therefore we use a simple discriminator architecture introduced by (Li et al.
-2020a) that contains very few parameters.
-To further address the risk of over-fitting, we perform differentiable data augmentation (Zhao et al.
-2020) on both the synthetic images output by the generator and the condensed samples before feeding them to the discriminator.
+[…]
 In deciding on the layer index at which the feature alignment will be employed, we considered the type of image features encoded by different parts of the teacher model.
@@
 Additionally, we posit that simply aligning the cumulative distribution of synthetic features from all classes with the real feature distribution is not ideal.
-Rather, we provide class-specific alignment using a conditional discriminator (Mirza and Osindero 2014).
+Rather, we provide class-specific alignment using a conditional discriminator (mirza2014conditional).
 The contrast between these two alternatives can be seen in the figure given in the appendix.
@@
 Formally, our real ($\mathcal{R}$) and fake ($\mathcal{F}$) sets can be defined as,
-=$\begin{split}\mathcal{R}&=\{(x,y)|(x,y)\in\mathcal{X}\}\\ \mathcal{F}&=\{(\hat{x},y_{ps})\}\cup\{(x,c)|(x,y)\in\mathcal{X},c\neq y\}\end{split}$ (4)
+$\begin{split}\mathcal{R}&=\{(x,y)|(x,y)\in\mathcal{X}\}\\ \mathcal{F}&=\{(\hat{x},y_{ps})\}\cup\{(x,c)|(x,y)\in\mathcal{X},c\neq y\}\end{split}$ (4)
 Combining condensed and synthetic samples:
@@
 However, as these methods can cause the student to be biased towards one data type, we avoided them.
-Instead, we extended the condensed dataset by adding the iteratively refined synthetic samples (through model inversion) and trained the student with randomly sampled batches from such union.
+Instead, we expanded the condensed dataset by adding the iteratively refined synthetic samples (through model inversion) and trained the student with randomly sampled batches from such union.
 Our distillation objective involves minimizing the distance between the predictions of the teacher and the student models, which can be summarized as,
@@
 Algorithm 1 Knowledge Distillation
-[…]
+INPUT: generator $G$, discriminatorDD, teacher $T$, student $S$ parameterized by $\theta_{S}$, condensed data $\mathcal{X}$
 OUTPUT: trained student $\theta^{*}_{S}$.
@@
 Experimental Evaluation
-To assess the effectiveness of our method, we incorporate condensed data guidance to the model inversion procedures of three state-of-the-art data-free KD methods and record the improvement.
+To assess the effectiveness of our method, we incorporate condensed data guidance to three state-of-the-art model-inversion methods and record the improvement in KD performance.
 These methods are Fast, CMI, and PRE-DFKD.
-Moreover, we also experiment with using limited real samples for feature regularization and observe the advantage against few-shot KD methods.
+Moreover, we also experiment with applying out method to expand limited real data and observe the advantage against few-shot KD methods.
 For this comparison, we selected NetGraft and FSKD as baselines.
@@
 To standardize the evaluation, we use fixed random seeds borrowed from the official implementations of the baselines.
-Datasets
-We use three image classification datasets, which are CIFAR-10/100 (Krizhevsky and Hinton 2009), and ImageNet-200 (Deng et al.
-2009).
-We conducted our experiments using condensed samples generated by three different methods, including those provided by (Zhao and Bilen 2021), (Cazenavette et al.
-2022) and (Zhao and Bilen 2023).
-While the results reported throughout our experiments primarily utilize the condensed samples from (Zhao and Bilen 2021), we also include an ablation study to compare the effectiveness of each condensation method.
-Implementation details
-We used the same generator architectures as shared in the official implementations of the data-free methods that we couple our method with (Fast, CMI, and PRE-DKD).
-Further details on generator and discriminator architectures as well as how we couple our method with individual MI methods can be found in the appendix.
-Condensed-data guided sample synthesis
-Tables 1 and 4 show the student accuracies upon coupling our approach with Fast, CMI, and PRE-DFKD.
-The results achieved by this coupling are indicated by the asterisk symbol with the annotation “*(ours w/ CS)”.
-To eliminate the advantage of having access to condensed samples, we also configure baselines where we combine these samples with the synthetic datasets generated by model inversion methods.
-These baselines are denoted with “+ CS” notations.
-Further, the row “Train w/ full real data” represents the accuracy of full-scale training of students on the target dataset and constitutes the upper bound.
-“Train w/ cond.
-samples” reflects the accuracy achieved by only using condensed samples for student training, which is the lower bound.
-In all experiments, we assume 50 spc (500 samples) and 10 spc (1000 samples) are available for CIFAR-10 and CIFAR-100, respectively.
-For ImageNet-200, we consider the availability of 10 spc (2000 samples).
-These represent 1%, 2%, and 2% of the total samples in their respective datasets.
 Dataset CIFAR-10 CIFAR-100 ImageNet-200
@@
 Teacher acc. 95.70 95.70 94.87 94.87 78.05 78.05 75.83 75.83 71.20 71.20
-Training student (S) with labeled data
+[…]
 Train w/ full real data
@@
 34.66 30.13 38.92 30.13 16.54 10.69 15.55 10.69 3.50 5.60
-Distilling student (S) with synthetic data
+[…]
 Fast 92.62 86.12 92.82 85.06 69.76 54.62 65.05 48.21 42.99 35.31
@@
 CMI* (ours w/ CS) 94.97 89.63 94.21 90.21 77.07 70.21 72.42 68.05 48.98 45.83
-Table 1: Student accuracies (%) obtained by enhancing Fast and CMI with condensed-sample guidance.
-Condensed datasets with 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200 were used.
+Table 1: Student accuracies (%) obtained by expanding condensed data using Fast and CMI. Condensed datasets with 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200 were used.
 ((a)) Synthetic feature distribution by data-free model inversion
@@
 Condensed data-guided synthetic samples exhibit better alignment with the real data distribution.
+Datasets
+We use three image classification datasets, which are CIFAR-10/100 (krizhevsky2009learning), and ImageNet-200 (deng2009imagenet).
+We conducted our experiments using condensed samples generated by three different methods, including those provided by (zhao2021siamese), (cazenavette2022dataset) and (zhao2023dataset).
+While the results reported throughout our experiments primarily utilize the condensed samples from (zhao2021siamese), we also include an ablation study to compare the effectiveness of each condensation method.
+Implementation details
+We used the same generator architectures as shared in the official implementations of the MI methods that we couple our method with (Fast, CMI, and PRE-DKD).
+Further details on generator and discriminator architectures as well as how we couple our method with individual MI methods can be found in the appendix.
+Impact of expanded condensed data on KD
+Tables 1 and 3 show the student accuracies upon coupling our approach with Fast, CMI, and PRE-DFKD.
+The results achieved by this coupling are indicated by the asterisk symbol with the annotation “*(ours w/ CS)”.
+We also configure naive baselines where we combine these samples with the synthetic datasets generated by model inversion methods.
+These are denoted with “+ CS” notations.
+Further, the row “Train w/ full real data” represents the accuracy of full-scale training of students on the target dataset and constitutes the upper bound.
+“Train w/ cond.
+samples” reflects the accuracy achieved by only using condensed samples for student training, which is the lower bound.
+In all experiments, we use condensed datasets of 50 samples per class (spc) (total 500 samples) and 10 spc (total 1000 samples) for CIFAR-10 and CIFAR-100, respectively.
+For ImageNet-200, we use 10 spc (total 2000 samples).
+These correspond to 1%, 2%, and 2% of the total samples in their respective datasets.
 First, we note that the performance of Fast method notably diminishes for pairs with low structural similarity (e.g.
-ResNet-34 (He et al.
-2016) & MobileNet-v2 (Howard et al.
-2017)).
-The improvement was especially significant for WRN-40-2 (Zagoruyko and Komodakis 2016b) & MobileNet-v2 pairs reaching up to $11.44\%$ on CIFAR-100.
+ResNet-34 (he2016deep) & MobileNet-v2 (howard2017mobilenets)).
+The improvement was especially significant for WRN-40-2 (zagoruyko2016wide) & MobileNet-v2 pairs reaching up to $11.44\%$ on CIFAR-100.
 CMI also has a considerable performance gap with respect to the upper bound for heterogeneous model pairs.
@@
 The results in Table 3 shows that our method effectively improves acuracy also for this baseline method.
-In all experiments, the simple inclusion of condensed samples in the distillation set (“+ CS”) neither mitigated this issue nor caused any substantial performance improvement in most cases.
-On the other hand, our condensed sample-guided model inversion (“*”) consistently increased student accuracy across different datasets and teacher-student pairs, which ensures that the benefit of our approach is not simply due to exposing the student to condensed samples during training.
+In all experiments, the simple combination of condensed samples and the synthetic samples from model inversion (“+ CS”) neither mitigated this issue nor caused any substantial performance improvement in most cases.
+On the other hand, our condensed sample-guided model inversion (“*”) consistently increased student accuracy across different datasets and teacher-student pairs, which ensures that the benefit of our approach is not simply due to exposing the student to more samples during training.
+Figure 4: Students distilled by only using samples from model inversion (left), and using expanded data generated by our method (right) for different condensed dataset set sizes.
+Condensed datasets with 1, 10, and 50 spc from CIFAR10, and 1 and 10 spc from CIFAR100, were used.
 Visual results
@@
 Figure 5: First two rows contain synthetic CIFAR-100 samples obtained with and w/o condensed data-guided model inversion. The last two show condensed and real samples.
-Figure 5: Students distilled by vanilla and condensed sample-guided model inversion for different condensed dataset set sizes.
-Condensed datasets with 1, 10, and 50 spc from CIFAR10, and 1 and 10 spc from CIFAR100, were used.
 Additionally, this conditioning does not compromise the diversity of the synthetic set, as demonstrated by the varied object views and scales in CMI*.
@@
 Thus, although the generated synthetic samples are visually realistic, they do not reveal any individual training samples.
-Moreover, to visually observe the effect of our method on the distribution of generated synthetic samples, we projected the high-dimensional feature space on a 2D plane using the t-SNE algorithm (Van der Maaten and Hinton 2008).
+Moreover, to visually observe the effect of our method on the distribution of generated synthetic samples, we projected the high-dimensional feature space on a 2D plane using the t-SNE algorithm (van2008visualizing).
 On the resulting plane, we compare the feature distributions observed for CMI and CMI*.
@@
 Each sample is marked with the color associated with its class label.
-We observe that synthetic samples generated by CMI are mostly clustered together and only partially correspond with real samples.
-Additionally, some feature clusters have almost no correspondence with real features (circled in red), meaning that the associated synthetic samples might not be contributing to the knowledge transfer.
+Synthetic samples generated by CMI are mostly clustered together and only partially correspond with real samples.
+Also, some feature clusters have almost no correspondence with real features (circled in red), meaning the associated synthetic samples might not be contributing to the knowledge transfer.
 In contrast, the synthetic feature distribution formed during CMI* shows better alignment with real data distribution.
 How does student accuracy scale with the condensed data size?
-After establishing that our proposed method can boost performance using condensed samples, we analyze how the scale of the available data affects the improvement.
+After establishing that our proposed method can boost the utility of condensed samples in KD, we analyze how the scale of the available data affects the improvement.
 For this, we consider condensed datasets of 3 different sizes (1 spc, 10 spc, 50 spc) for CIFAR-10 and 2 different sizes (1 spc, 10 spc) for CIFAR-100.
@@
 Does the condensation method affect the quality of model inversion?
-We study the impact of the dataset condensation method used to produce the condensed samples on the effectiveness of our approach and display the results in Table 2.
-DSA, DM and MTT refer to (Zhao and Bilen 2021), (Zhao and Bilen 2023), and (Cazenavette et al.
-2022) respectively.
-The results indicate that while the choice of condensation method can impact the final student accuracy, our method consistently improves performance across all types of condensed samples tested.
-This suggests that its effectiveness is not dependent on any single condensation approach.
+We study the impact of the dataset condensation used to produce the condensed samples on the effectiveness of our approach and display the results in Table 2.
 Dataset CIFAR-10
@@
 Student MobileNet-v2 MobileNet-v2
-Ablation Study on Different Dataset Condensation Methods
+[…]
 KD Method CS Type DSA DM MTT DSA DM MTT
@@
 Fast* (ours w/ CS) 88.05 87.82 87.11 89.24 90.30 90.38
-Table 2: Impact of different types of condensed samples (CS) on student accuracy (%) for different dataset condensation strategies. Condensed dataset with 50 spc from CIFAR10 was used.
+Table 2: Impact of different types of condensed samples (CS) on student accuracy (%) for different dataset condensation strategies. Dataset contains 50 spc from CIFAR10.
+DSA, DM and MTT refer to (zhao2021siamese), (zhao2023dataset), and (cazenavette2022dataset) respectively.
+The results indicate that while the choice of condensation method can impact the final student accuracy, our method consistently improves performance across all types of condensed samples tested.
+This suggests that its effectiveness is not dependent on any single condensation approach.
+Dataset CIFAR-10 CIFAR-100 ImageNet-200
+Teacher ResNet-34 WRN-40-2 ResNet-34 WRN-40-2 ResNet-34 ResNet-34
+Student MobileNet-v2 MobileNet-v2 MobileNet-v2 MobileNet-v2 ResNet-18 MobileNet-v2
+Teacher acc. 95.70 94.87 78.05 75.83 71.20 71.20
+[…]
+Train w/ full real data 93.79 93.79 72.80 72.80 64.90 55.06
+Train w/ cond. samples (CS) 27.24 27.24 10.69 10.69 3.50 5.60
+Train w/ few real samples (RS) 27.55 27.55 4.66 4.66 1.42 1.30
+[…]
+PRE-DFKD 83.12 83.12 66.56 61.83 54.20 47.26
+PRE-DFKD* (ours w/ CS) 86.89 88.34 67.39 63.41 55.22 49.73
+PRE-DFKD* (ours w/ RS) 86.77 87.47 71.63 64.37 54.95 49.68
+Table 3: Student accuracies (%) obtained by utilizing the expanded condensed and real datasets set using PRE-DFKD (CS and RS).
+Condensed and real sample sets contain 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200.
+[…]
+Table 4: Comparison with different few-shot baselines using few real samples.
 Ablation study
@@
 + class-specific alignment (CMI*) 94.21
-Table 3: MobileNet-v2 student accuracy on CIFAR10 for various configurations of the components used in our method. Condensed dataset with 50 spc from CIFAR10 was used.
-Dataset CIFAR-10 CIFAR-100 ImageNet-200
-Teacher ResNet-34 WRN-40-2 ResNet-34 WRN-40-2 ResNet-34 ResNet-34
-Student MobileNet-v2 MobileNet-v2 MobileNet-v2 MobileNet-v2 ResNet-18 MobileNet-v2
-Teacher acc. 95.70 94.87 78.05 75.83 71.20 71.20
-Accuracy of student (S) trained with labeled data
-Train w/ full real data 93.79 93.79 72.80 72.80 64.90 55.06
-Train w/ cond. samples (CS) 27.24 27.24 10.69 10.69 3.50 5.60
-Train w/ few real samples (RS) 27.55 27.55 4.66 4.66 1.42 1.30
-Accuracy of student (S) trained with synthetic data
-PRE-DFKD 83.12 83.12 66.56 61.83 54.20 47.26
-PRE-DFKD* (ours w/ CS) 86.89 88.34 67.39 63.41 55.22 49.73
-PRE-DFKD* (ours w/ RS) 86.77 87.47 71.63 64.37 54.95 49.68
-Table 4: Student accuracies (%) obtained by enhancing PRE-DFKD with condensed and real-sample guidance (CS and RS).
-Condensed and real sample sets containing 50 spc from CIFAR10, 10 spc from CIFAR100, and 10 spc from ImageNet-200 were used.
-Dataset CIFAR-10 CIFAR100 Teacher VGG-11 VGG-11 VGG-11 VGG-11 Student VGG-11 (50% pruned) VGG-11 (75% pruned) VGG-11 (25% pruned) VGG-11 (50% pruned) Teacher acc.
-92.25 92.25 71.23 71.23 Few-shot distillation accuracy FSKD 78.69 36.00 54.72 24.73 PRE-DFKD* 83.26 68.48 63.81 61.21 ((a)) Student accuracy comparison with FSKD baseline.
-Real sample sets containing 50 spc from CIFAR10, and 10 spc from CIFAR100 were used.
-Dataset CIFAR-10 CIFAR100 Teacher VGG-16 VGG-16 Student ResNet-18 ResNet-18 Teacher acc.
-94.16 74.00 Few-shot distillation accuracy NetGraft 73.69 55.51 PRE-DFKD* 88.38 68.54 ((b)) Comparison with few-shot NetGraft baseline.
-1 spc from both CIFAR10 / 100 were used.
-Table 5: Comparison with different few-shot baselines using few real samples.
-This study involves student accuracies upon enhancing a base model inversion method (CMI) in 3 different settings.
+Table 5: MobileNet-v2 student accuracy on CIFAR10 for various configurations of the components used in our method. Dataset with 50 spc from CIFAR10 was used.
+This involves student accuracies upon enhancing a base model inversion method (CMI) in 3 different settings.
 (1) utilizing condensed samples only by appending them to the distillation set (“+ CS”), (2) using them also to regularize the feature distribution of synthetic data (“+ CS-guided MI”) (3) the final version of our condensed sample-guided model inversion (CMI*).
@@
 Later, we include the final component, i.e., class-specific alignment, and record student accuracy (“+ class-specific alignment”).
-Although this addition caused only a marginal increase in performance, we still retain it as it improves results without introducing any notable computational burden.
-Real-data guided sample synthesis
-Our method is also inherently capable of utilizing limited real samples from the target dataset when provided.
+This caused only a marginal increase in performance but we retain it as it improves results without introducing any notable computational burden.
+Impact of expanded real data on KD
+Our method is also inherently capable of expanding limited real samples from the target dataset in few-shot scenarios.
 To show this, we repeated the experiments in Table 3 by replacing the condensed samples with real ones.
@@
 The overall improvement yielded by our method upon utilizing few real samples is comparable that using condensed ones.
-This highlights the effectiveness of our approach in few-shot KD scenarios.
 Comparison with few-shot methods
@@
 As for the comparison with Netgraft, we use VGG16 teachers and ResNet-18 students, with 1 spc from both datasets.
-From Table 5, it can be observed that FSKD is only effective for low pruning rates while performing poorly for higher rates.
+From Table 4, it can be observed FSKD is only effective for low pruning rates while performing poorly for higher rates.
 Similarly, the accuracies achieved by NetGraft were also much lower than the upper-bound as seen in Table 4.
@@
 Conclusion
-In conclusion, our study addresses the challenge of enhancing data-free knowledge distillation by incorporating supplementary information available from the target dataset—something traditional model inversion methods fail to leverage.
-Recognizing the potential of condensed samples to serve as this supplementary information, we developed a method that refines the synthetic data generation process within the model inversion framework.
-This integration helps align the synthetic data more closely with the real data distribution, effectively reducing the domain gap that often undermines KD performance.
-Our experiments demonstrate the effectiveness of this approach, showing accuracy improvements of up to $11.4\%$ across various datasets and model configurations.
-The method is also applicable in environments where only a few real samples are accessible, such as in few-shot learning scenarios.
-Acknowledgement
-This work is supported by the National Research Foundation, Singapore under its Competitive Research Programme Award NRF-CRP23-2019-0003 and Singapore Ministry of Education Academic Research Fund T1 251RES1905.
-We thank Cihan Acar and Liu Siying for helpful comments on the manuscript.
-Appendix A: Computing Infrastructure
-The specifications of the system used to run computational experiments are:
-  • Ubuntu 22.04.4 LTS
-  • NVIDIA A100 80GB GPUs
-  • Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz
-Appendix B: Implementation Details
-Generator architectures
-The Fast and CMI baselines share the same generator architecture, which is detailed in Table 6.
-The PRE-DFKD generator differs slightly, with the inclusion of spectral normalization following the convolution operations, as outlined in Table 7.
-In addition, the final activation function is hyperbolic tangent (Tanh), as opposed to sigmoid.
-Both tables utilize ”$nz$” to represent the dimension of the random vector that is drawn from a standard normal distribution.
-In PRE-DFKD, the value of ”$nz$” is set to 1000, while in Fast and CMI, it is set to 256.
-Lastly, the described architectures are cofigured to produce 32x32 RGB images (for CIFAR-10 and CIFAR-100).
-The activation shapes vary for datasets with samples of higher resolution (e.g.
-ImageNet-200).
-Layer # Operation in shape out shape
-1 Linear nz 8x8x128
-2 BN 8x8x128 8x8x128
-3 2xUpsample 8x8x128 16x16x128
-4 Conv-BN-ReLU 16x16x128 16x16x128
-5 2xUpsample 16x16x128 32x32x128
-6 Conv-BN-ReLU 32x32x128 32x32x64
-7 Conv 32x32x64 32x32x3
-8 Sigmoid 32x32x3 32x32x3
-Table 6: Generator architecture used in Fast and CMI experiments
-Layer # Operation in shape out shape
-1 Linear nz 8x8x128
-2 BN 8x8x128 8x8x128
-3 2xUpsample 8x8x128 16x16x128
-4 Conv-SN-BN-ReLU 16x16x128 16x16x128
-5 2xUpsample 16x16x128 32x32x128
-6 Conv-SN-BN-ReLU 32x32x128 32x32x64
-7 Conv-SN 32x32x64 32x32x3
-8 Tanh 32x32x3 32x32x3
-9 BN 32x32x3 32x32x3
-Table 7: Generator architecture used in PRE-DFKD experiments. SN stands for spectral normalization.
-Discriminator architecture & training details
-For all experiments, we utilize the same discriminator whose architecture is described in Table 3.
-“nz” has the same meaning as in Tables 6 and 7.
-“nc” corresponds to the number of classes, for instance “nc=100” for CIFAR-100 dataset.
-To update the discriminator we use Adam optimizer with learning rate=$0.002$ and beta values=$[0.5,0.999]$.
-While propagating the discriminator loss to the generator, we scale it with the coefficient of $0.5$, thus the generator loss $L_{G}$ in Equation 2 becomes:
-${L}_{G}=\mathcal{L}_{MI}+0.5\times\mathcal{L}_{FA}$ (6)
-Layer # Operation in shape out shape
-1 Linear nz+nc 1024
-2 ReLU 1024 1024
-3 Dropout (0.5) 1024 1024
-4 Linear 1024 1
-Table 8: Architectural details of the feature discriminator.
-Coupling with model inversion methods
-Fast and CMI use only a single generator, so we simply attach our discriminator as shown in Figure 1(a) in the main paper.
-In addition, we notice that both methods reset the generator periodically by randomly re-initializing its weights.
-To prevent any instability in our adversarial feature regularization setup, we also reset the discriminator with the same frequency.
-We employ the Fast-2 configuration on CIFAR-10 and CIFAR-100 datasets, which update the generator two times for every update that the student receives.
-Meanwhile, for ImageNet-200, we adhere to the Fast-50 configuration, which is consistent with the original paper’s approach for complex datasets.
-As the PRE-DFKD framework contains 2 generators, namely “novel” and “memory” generators, attaching our method requires further consideration.
-While the “novel” generator is responsible for producing new/informative samples in the framework, the memory generator memorizes the cumulative output distribution of the novel generator.
-As conditioning the novel generator also implicitly affects the memory generator, we only pair the discriminator with it.
-We run CMI, Fast, and PRE-DFKD for 200, 220, and 220 epochs, respectively.
-All results reflect the maximum student accuracy observed during distillation.
-Moreover, PRE-DFKD does not store generated samples and instead distills the student on-the-fly with newly generated samples at each epoch.
-This prevents us from directly applying Algorithm 1 to combine condensed and synthetic samples, and the other alternatives discussed in Section 3 impair student accuracy.
-For this reason, we only use the condensed samples to guide synthetic data generation and exclude them while training the student.
-Even without utilizing condensed samples in a direct way in distillation, we observe accuracy improvement on both datasets.
-Appendix C: Condensed Datasets
-The basic dataset condensation process uses a reference model and optimizes condensed samples to match the gradients that larger real batches produce for that model.
-In our experiments, we used condensed samples retrieved from the official GitHub pages of (Zhao, Mopuri, and Bilen 2021) and (Cazenavette et al.
-2022).
-Both works, used the same 3-layer CNN architecture as the reference model to generate these datasets.
-For CIFAR-10, we used samples generated by both (Zhao and Bilen 2021) and (Zhao and Bilen 2023).
-The samples from (Zhao and Bilen 2021) are available in three sizes: 1, 10, and 50 samples per class (spc), while (Zhao and Bilen 2023) provides samples only at 50 spc.
-For CIFAR-100, only the 1 and 10 spc settings from (Zhao and Bilen 2021) are available.
-For ImageNet-200, we utilized the condensed samples from (Cazenavette et al.
-2022), which provides 10 spc.
-It is worth noting that none of the teacher models in our experiments were used as reference models during the generation of these samples.
-This demonstrates that our method can effectively utilize condensed datasets, regardless of the reference model used in their optimization process.
-Appendix D: Time complexity of model inversion
-To show that our method does not sacrifice significant speed while enhancing the performance, we share the model inversion times in table 9.
-Method Time (hrs)
-Fast 3
-Fast* 4
-CMI 11
-CMI* 13.5
-PRE-DFKD 11.6
-PRE-DFKD* 13.5
-Table 9: Times consumed by different model inversion methods.
-Appendix E: Class-specific Feature Alignment
-When the discriminator is configured to only differentiate between the generated synthetic samples and the condensed ones, the generator can achieve its optimization objective by merely aligning the two distributions, as demonstrated in Figure 6(a).
-However, for optimal results, it is preferable to align the synthetic samples with the condensed ones based on their respective classes, as shown in Figure 6(b).
-For instance, a synthetic car image should exhibit similar semantic features to a condensed car image rather than a condensed sunflower image.
-((a)) generic alignment
-((b)) class-specific alignment
-Figure 6: Illustration of feature spaces observed after different feature alignment strategies.
-Condensed (C) and synthetic (S) features are denoted by red and blue colors respectively.
-Each shape indicates a unique class ID (e.g.
-c1).
+We address the challenge of limited information in condensed datasets, which often hinders their effectiveness in KD.
+We propose a method that leverages MI to expand these condensed datasets, generating synthetic data that enriches the compact representation and improves KD performance.
+Our approach utilizes condensed samples as prototypes to guide the MI process, ensuring that the generated synthetic data aligns closely with the underlying data distribution represented by the condensed set.
+Our experiments demonstrate the effectiveness of our method across various datasets, model architectures, and state-of-the-art MI techniques (Fast, CMI, PRE-DFKD).
+The results show consistent improvements in KD, compared to using condensed datasets alone or standard MI-based KD methods.
+Our approach also applies to few-shot learning, outperforming existing few-shot KD by effectively leveraging limited real data for synthetic data generation.